OpenStudy (anonymous):

Find the product. (3p^4)^3 · (p^2)^7

OpenStudy (whpalmer4):

This is very much like the last one we did. What would you like to do first?

OpenStudy (whpalmer4):

Why don't you simplify the \((p^2)^7\) part first — what does that give you?

OpenStudy (anonymous):

p^14

OpenStudy (whpalmer4):

Yep. How about the first part?

OpenStudy (whpalmer4):

I would split it apart like we did last time...

OpenStudy (anonymous):

So... (3^4)^3... Which is 3^12?

OpenStudy (whpalmer4):

Take another look, isn't it \((3p^4)^3\) ?

OpenStudy (whpalmer4):

Remember, \[(ab)^n = a^nb^n\]so we can split it into \[(3)^3(p^4)^3\]

OpenStudy (anonymous):

Oh! Okay. So... 9p^12?

OpenStudy (whpalmer4):

Have to keep the exponents together with the bases they came with, no date-swapping here :-) Getting closer...the \(p^{12}\) part is correct, but the number is incorrect.

OpenStudy (whpalmer4):

\[(3)^1 = 3\]\[(3)^2 = 3*3 = 9\]\[(3)^3 = 3*3*3 = \]

OpenStudy (anonymous):

Ohhh! Sorry. I did 3*3... Not 3*3*3. So it's 27p^12.

OpenStudy (whpalmer4):

Right. Now you have to multiply \(27p^{12} p^{14}\)

OpenStudy (anonymous):

So, 27p^26?

OpenStudy (whpalmer4):

Correctamundo!

OpenStudy (whpalmer4):

Got some more of these that you can use to impress me? :-)

OpenStudy (anonymous):

Haha I'll message you some questions and answers I come up with so you can tell me if I'm getting them right :)

OpenStudy (whpalmer4):

Sounds like a plan!

OpenStudy (anonymous):

Again, Thank you! :)

OpenStudy (whpalmer4):

You're welcome, of course!